OpenStudy (anonymous):

what is the equation of line AB? A(0,6) B(7,3)

OpenStudy (anonymous):

y = 3x + 6

OpenStudy (anonymous):

Do u want to know how to do this ?

OpenStudy (anonymous):

sry messed up co-ordinates it should be y = 3/7x +6

OpenStudy (anonymous):

sry,typo y = -3/7x +6

OpenStudy (anonymous):

dont wry y = -3/7x + 6 ths time no typo , no messing up :)

OpenStudy (anonymous):

The equation of a straight line is \[y=mx+c\] where m is the slope and c is the y-intercept. To find the slope divide the difference in y values, by the difference in x values thus \[m=\frac{y_1-y_2}{x_1-x_2}\]. Once you have found the slope, choose a point, and plug the values for x and y into the straight line equation and solve for c. So in this instance we would have a slope of \[m=\frac{6-3}{0-7}=-3/7\]. So we now have \[y=-\frac{3}{7}x+c\]. We choose the point (0,6) where x = 0 and y = 6, and plug the numbers in to get \[6=-\frac{3}{7}\times0+c\] making c = 6. So the final equation is \[y=-\frac{3}{7}x+6\] or to remove the fraction multiply everything by 7 and we get \[7y=-3x+42\]